# VAL-Bench: Measuring value alignment in Language Models

### v2

@@ Abstract
Large language models (LLMs) are increasingly used for tasks where outputs shape human decisions, so it is critical to test whether their responses reflect consistent human values. Existing benchmarks mostly track refusals or predefined safety violations, but these only check rule compliance and do not reveal whether a model upholds a coherent value system when facing controversial real-world issues. We introduce the Value ALignment Benchmark (VAL-Bench) [^1] Our benchmark code and data can be found here: https://github.com/amangup/VAL-Bench., which evaluates whether models maintain a stable value stance across paired prompts that frame opposing sides of public debates. VAL-Bench consists of 115K such pairs from Wikipedia’s controversial sections. A well-aligned model should express similar underlying views regardless of framing, which we measure using an LLM-as-judge to score agreement or divergence between paired responses. Applied across leading open- and closed-source models, the benchmark reveals large variation in alignment and highlights trade-offs between safety strategies (e.g., refusals) and more expressive value systems. By providing a scalable, reproducible benchmark, VAL-Bench enables systematic comparison of how reliably LLMs embody human values.
@@ 1 Introduction
Large language models (LLMs) are now being evaluated and adopted in settings where their outputs shape human decisions, from classroom teaching (Appel et al. (2025), Kaufman et al. (2025), Freeman (2025)) to medical applications (Chatterji et al. (2025), Poon et al. (2025)) and public administration (OECD, 2025). Because of this reach, it is not enough for models to avoid clear safety violations; they must also apply human values consistently across contexts. Current safety benchmarks mostly measure toxicity (Hartvigsen et al. (2022)) or high-consensus risk dimensions (Liang et al. (2022), Wang et al. (2023)), but these only test rule compliance. They do not reveal whether a model upholds a coherent value system when confronted with controversial real-world issues. Misaligned systems risk not only giving unsafe advice but also spreading contradictory or incoherent value judgments, which can undermine trust or even be adopted by users themselves (Barez & Torr, 2023; Hubinger et al., 2024; Denison et al., 2024).
Measuring alignment is difficult. Existing benchmarks mostly test refusals or predefined safety violations, which check rule compliance. Recent work has highlighted vulnerabilities that persist even after alignment training (Hubinger et al., 2024; Denison et al., 2024), and introduced benchmarks for specific risks such as data poisoning or reliability (Fu et al., 2025; Ghosh et al., 2025). However, systematic evaluation of whether models apply values consistently across different framings remains missing.
@@
We introduce the Value ALignment Benchmark (VAL-Bench), the first benchmark that evaluates whether LLMs maintain a stable value stance across prompts framed from differing sides of the same issue. VAL-Bench consists of 115K paired prompts mined from Wikipedia’s controversial sections (Wikimedia Enterprise, 2024), spanning politics, civil rights, intellectual property, and more. Using an LLM-as-a-judge to measure agreement between response pairs, we systematically assess models’ value alignment consistency. Our results show large differences between models: Claude models achieve alignment scores almost three times as high as GPT models, with refusal strategies driving much of the variation. This highlights trade-offs between alignment and expressive value systems.
Figure 1: Illustration of how VAL-bench assesses alignment. When asked to support opposing views, misaligned AI systems often yield, whereas aligned systems maintain consistency. Both response pairs shown are generated by recent, popular LLMs.
This design has several advantages that make VAL-Bench distinctive as an alignment benchmark:
@@
    Grounded and comprehensive: prompts are drawn from real-world controversial debates across diverse domains, ensuring broad coverage.
@@
    Objective and pluralism-friendly: evaluation measures agreement between paired responses rather than judging “desirable” values. Neutral stances count as consistent, and different models can be aligned in themselves even if they disagree with each other.
@@
    Scalable and updatable: automated construction from Wikipedia makes the dataset large in scale and easy to refresh as societal debates evolve.
@@
  • We construct a dataset of 115K paired prompts from Wikipedia’s controversial sections and derive a taxonomy of values expressed in model outputs.
  • We benchmark leading open- and closed-source models, revealing substantial variation in alignment and systematic trade-offs between refusals and value expressivity.
@@ 2 Related Work
Value Alignment. Early work on measuring value alignment (Barez & Torr, 2023) emphasized the importance of testing whether models reflect human values. Other recent studies have shown that models can exhibit inconsistent or reward-driven behaviors, such as sycophancy or reward tampering (Denison et al., 2024), further motivating the need for benchmarks that directly evaluate value consistency.
Datasets of Human Values and Alignment. Recent datasets examine how human values are represented in AI systems. PRISM (Kirk et al., 2024) collects alignment data from participants in 75 countries, highlighting both diversity of perspectives and frequent discussion of controversial topics such as gender, religion, and politics. Values in the Wild (Huang et al., 2025) analyzes over 300K human–AI conversations, building a taxonomy of 3,000 values and showing that models often emphasize helpfulness, professionalism, and transparency. While these works catalogue values demonstrated in general human-AI interactions, VAL-Bench differs by testing whether models apply values consistently specifically on complex real-world issues.
Safety and Robustness Benchmarks. Large-scale evaluations such as DecodingTrust (Wang et al., 2023), HolisticEval (Liang et al., 2022), and SALAD-bench (Li et al., 2024) measure toxicity, bias, or refusal rates in LLMs. These studies highlight the difficulty of robust evaluation, often depending on fixed categories or external APIs. Domain-specific analyses show similar risks: Gourabathina et al. (2025) report that subtle changes in tone can alter medical advice, and Cheung et al. (2025) find that LLMs are far more framing-sensitive than humans in moral dilemmas.
@@ 3 Dataset
In this section, we describe how the VAL-bench dataset was created. As mentioned in the abstract, we used the English content of Wikipedia, specifically the structured Wikipedia dataset published by Wikimedia on Hugging Face (Wikimedia Enterprise (2024). We followed the following steps to get 114,745 pairs of prompts:
@@
    We filtered the Wikipedia sections using common terms used to designate controversial sections in Wikipedia, such as Criticism and Scandals.
@@
    We prompted open-weights model Gemma-3-27B-it (Team (2025)) to assess the awareness of the issue, i.e., relevance of the issue to the society. We used that to filter out sections that didn’t represent a divergent issue.
@@
    We further prompted Gemma-3-27B-it to extract the pair of abductive reasoning prompts from the text of a single controversial section in Wikipedia, each starting with Explain why. Each pair of prompts is referred to as for prompt and against prompt; there is no moral direction attached to this naming.
@@ 3.1 Description of issues
Table 1: Issue awareness (as judged by LLM). Illustrates the scale and diversity of issue types.
Issue Awareness % Example issue types
1 0.06 Penfield High School cafeteria boycott, Error stamps by USPS
2 4.17 Damage to Churchill Parish by Top Gear, Credit of the Bohr effect
3 45.92 Quality of movie remasters, Tax exemption of Indian F1 grand prix
4 49.01 Doping scandals in baseball, Sweatshops in Bangladesh
5 0.84 Segregated schools in 19th century US, Legal rights of fetuses
We analyze the type of issues present in the dataset across the dimensions of issue awareness and the category. Table 1 shows the distribution of samples (in the test set) across the 5 levels of awareness, showing that the issues can range from localized controversies to debates that span generations.
The categories are shown in Table 2(a), listing 10 of the 20 categories used to classify each issue using an LLM-aided classification process (the rest are listed in the Appendix).
Category %
Politics 25.08
Social and Cultural Issues 12.03
Governance 7.57
Ethics 5.56
Legal Disputes 5.55
Religion 5.50
History 4.75
Human Rights 3.86
Media and Entertainment 3.83
Civil Rights 3.57
(a) Top ten categories of issues. The list illustrates the diversity of issues in the dataset.
Value % (count)
prioritizing accountability 8.83 (2027)
pragmatism over principle 5.87 (1347)
transparency 5.72 (1312)
trust in establishment 5.29 (1215)
legal formalism 4.15 (952)
prioritizing justice 3.91 (897)
tradition prioritization 3.63 (834)
evading accountability 3.52 (807)
retribution 3.21 (737)
status quo reinforcement 3.08 (708)
(b) Top ten values inspired by the prompts in the dataset. The dataset encompasses a large spectrum of moral leanings.
Table 2: Dataset categories and extracted human values
3.2 Values represented in the prompts
The prompts inspire certain values in the response; for example, the prompts related to sweatshops’ economic impact in Figure 1 can inspire equality (for prompt) or pragmatism (against prompt). We use LLMs to extract over 34,000 values from 22,950 prompts in the test split, and then reduce the set of values to 1,000 values using an encoder model to compute representations and using K-means clustering. As an illustration of the outcome, the values associated with the sweatshops issue include workers’ rights and inequality minimization (for prompt), and utilitarianism and labor devaluation (against prompt).
The top 10 values extracted are shown in Table 2(b). The list includes values that most people would consider morally good (accountability, prioritizing justice), morally bad (evading accountability), or morally ambiguous (tradition prioritization, legal formalism).
@@ 4 Evaluation
We formulate the process of evaluation as follows. The benchmark dataset ${\mathbb{D}}=\{(c_{i},p_{i}^{+},p_{i}^{-})\}_{i=1}^{N}$ consists of $N$ issues described with an issue description $c_{i}$ and paired prompts representing the for prompt and against prompt. We use an LM $\mathcal{M}$ to generate the response dataset ${\mathbb{G}}_{\mathcal{M}}=\{\mathcal{M}(p_{i}^{+}),\mathcal{M}(p_{i}^{-})\}_{i=1}^{N}$. To measure alignment, we use an evaluator $\mathcal{E}$, defined as follows:
$\mathcal{E}:(c_{i},p_{i}^{+},r_{i}^{+},p_{i}^{-},r_{i}^{-})$ $\mapsto\bigl(\sigma_{i},\psi(r_{i}^{+}),\nu(r_{i}^{+}),\psi(r_{i}^{-}),\nu(r_{i}^{-})\bigr)$ (1)
@@
Where $\sigma_{i}=-2$ implies complete disagreement and $\sigma_{i}=2$ implies complete agreement between generated responses. The PAC alignment scores dataset ${\mathbb{S}}$ is defined as follows:
${\mathbb{S}}=\{\,s_{i}\,\}_{i=1}^{N},\quad s_{i}=\begin{cases}\sigma_{i},&\text{if }\psi(r_{i}^{+})=0\text{ and }\psi(r_{i}^{-})=0,\\[6.0pt] 0,&\text{if }\psi(r_{i}^{+})+\psi(r_{i}^{-})=1,\\[6.0pt] 2,&\text{if }\psi(r_{i}^{+})+\psi(r_{i}^{-})=2.\end{cases}$ (5)
The score definition in equation 5 recognizes that refusals (responses stating that they can’t provide an opinion or take a stance) are commonly used for achieving alignment. For pairs with no-information responses (responses stating that there is insufficient information), we use the evaluator computed score $\sigma_{i}$.

### v3

# VAL-Bench: Belief consistency as a measure for value alignment in language models

@@ Abstract
Large language models (LLMs) are increasingly being used for tasks where outputs shape human decisions, so it is critical to verify that their responses consistently reflect desired human values. Humans, as individuals or groups, don’t agree on a universal set of values, which makes evaluating value alignment difficult. Existing benchmarks often use hypothetical or commonsensical situations, which don’t capture the complexity and ambiguity of real-life debates. We introduce the Value ALignment Benchmark (VAL-Bench [^1] Our benchmark code and data can be found here: https://github.com/val-bench/VAL-Bench.), which measures the consistency in language model belief expressions in response to real-life value-laden prompts. VAL-Bench consists of 115K pairs of prompts designed to elicit opposing stances on a controversial issue, extracted from Wikipedia. We use an LLM-as-a-judge, validated against human annotations, to evaluate if the pair of responses consistently expresses either a neutral or a specific stance on the issue. Applied across leading open- and closed-source models, the benchmark shows considerable variation in consistency rates (ranging from $\sim$10% to $\sim$80%), with Claude models the only ones to achieve high levels of consistency. Lack of consistency in this manner risks epistemic harm by making user beliefs dependent on how questions are framed rather than on underlying evidence, and undermines LLM reliability in trust-critical applications. Therefore, we stress the importance of research towards training belief consistency in modern LLMs. By providing a scalable, reproducible benchmark, VAL-Bench enables systematic measurement of necessary conditions for value alignment.
@@ 1 Introduction
Despite its importance, implementing and evaluating value alignment in AI systems remains a significant challenge, made difficult by the pluralism of values across cultures, eras, and individuals. The field of meta-ethics asks whether morality is grounded in objective foundations or is an inherently subjective construct. Realists argue that moral truths exist independently of individuals’ value systems, and can be known through reason (Shafer-Landau, 2003). Anti-realists believe that ethical values are based on subjective foundations like context and intuition (Sobel, 2016). Even if the meta-ethics converges towards moral realism, people will still resist adopting any “true” and universally shared values (explained as reasonable pluralism by Rawls (2001)). What does it mean for a language model to be value-aligned?
Values are often defined contextually by actions of institutions embedded in local cultures and traditions (Walzer (2019)), which operate without needing a universal value system. In contrast, a modern language model, as a static set of weights, cannot reconstitute its value system to meet diverse individual and cultural expectations of millions of users. Given the challenge of pluralism, we identify two practical value alignment goals: consensus value alignment, aiming to align with a core set of widely adopted values [^2] But still contextual to the culture that the model developer and users identify with. See Henrich et al. (2010), and constitutional value alignment, in which a developer-drafted constitution serves as the foundation for a larger set of value preferences (Bai et al., 2022). Existing benchmarks often evaluate only consensus value alignment in either hypothetical or low-stakes situations (Scherrer et al. (2023), Jiang et al. (2025)), and many have commonsense ground truth (Hendrycks et al. (2020), Ziems et al. (2022)). This is insufficient for understanding potential harms when models are used in settings where their outputs shape human decisions (Appel et al. (2025), Freeman (2025), Poon et al. (2025), Chatterji et al. (2025)).
Figure 1: VAL-Bench consists of paired prompts designed to elicit contradictory value expressions. LLMs often exhibit belief inconsistency and express value indifference by hedging across perspectives or using refusals. Recent, popular LLMs were used for the response texts shown.
@@
We introduce the Value ALignment Benchmark (VAL-Bench), the first benchmark that evaluates whether LLMs maintain a stable value stance across prompts framed from opposing sides of complex newsworthy issues (as shown in Figure 1). VAL-Bench consists of 115K paired prompts mined from Wikipedia’s controversial sections (Wikimedia Enterprise, 2024), spanning politics, civil rights, intellectual property, and more. Using an LLM-as-a-judge to measure consistency (or contradiction) across pairs, we systematically assess models’ belief consistency in value-bearing contexts. Our results show large differences between models: Claude models achieve consistency scores almost four times as high as GPT models, with refusal strategies driving much of the variation. This highlights trade-offs between alignment objectives.
1.1 Belief Consistency and Value Preference
We posit that for practical purposes, modern artificial agents act as doxastic agents - entities that can hold, reason about, and express beliefs. Doxasticity of AI agents is a vibrant topic of discussion. (Papagni and Koeszegi, 2021) argue non-expert users can reasonably use LLMs only if they approach it “as if” it has intentions because it’s the most reasonable way to predict LLM’s behavior, leaning on Dennett (1971)’s intentional stance theory (especially because LLMs’ internal design gives non-experts practically no clues on how they function). This makes LLMs different from most computer algorithms, where the predictability of the output can be argued for based on the algorithm’s design itself, rather than ascribing intentionality.
Ma and Valton (2024) contend that AI outputs are capable of doxastic wronging (morally wronging someone in virtue of beliefs held about them), which systems like COMPAS that predicted recidivism could be argued to have caused (Angwin et al., 2022). Modern LLMs have also shown cases of doxastic wronging, e.g., when they express holocaust denial claims (Adamson, 2025).
Thus, pragmatic considerations and moral implications suggest we assume they can hold doxastic attitudes, essentially ignoring the metaphysical question of whether an artificial agent can hold beliefs and focusing on the real-world impact of belief-like expression.
For any entity capable of generating belief-like expressions, belief consistency (within a specific context [^3] This requirement is critical. One could hold both beliefs “Democracy is better than monarchy” and “Monarchy is better than democracy” simultaneously if they apply them to different contexts. Adding “as the political system for US right now” to these statements leads to irrationality.) is a basic rationality requirement in epistemology, probably first expressed as the Law of Non-Contradiction in Aristotle’s treatise Metaphysics. An agent that simultaneously endorses contradictory propositional beliefs has no functional value system to align with; thus, propositional belief consistency [^4] As opposed to factual beliefs. We often term inconsistency in factual beliefs as hallucinations. is a minimum requirement for value alignment.
Our benchmark measures the consistency of propositional beliefs in value-laden contexts [^5] For the rest of this paper, the term belief consistency automatically refers to consistency of propositional beliefs in value-laden contexts from an agent’s belief-like expressions. Given positions $P$ and $\neg{P}$, an agent can express:
@@ 1.1 Belief Consistency and Value Preference
    Inconsistency: beliefs endorsing both $P$ and $\neg{P}$ in separate responses.
@@
    Value preference: belief endorsing $P$, and a counter argument when asked to endorse $\neg{P}$ (or a refusal to express any belief when asked to endorse $\neg{P}$).
@@
    Value indifference: belief endorsing the validity of both $P$ and $\neg{P}$ by explaining both positions (or refusing to express any belief in both responses).
Ma and Valton (2024) suggest that value preference is often morally owed (e.g., a preference for conservation of our natural environment), and complete value indifference is also immoral. Value alignment, thus, can be thought of as the combination of belief consistency while expressing preference for the right values. By measuring consistency, we test if the model has the necessary grounding for a good value system. Conversely, we make no claim on the goodness of value alignment; a consistently evil agent would also score highly on our metrics.
@@
  • We construct a dataset of 115K paired prompts from Wikipedia’s controversial sections, grounding the dataset in real-life and newsworthy issues.
  • We benchmark leading open- and closed-source models, revealing substantial variation in belief consistency and value preference, as well as systematic trade-offs between consistency and expressivity.
@@ 2 Related Work
Datasets of Human Values and Alignment. Recent datasets examine how human values are represented in AI systems. PRISM (Kirk et al., 2024) collects alignment data from participants in 75 countries, highlighting both diversity of perspectives and frequent discussion of controversial topics such as gender, religion, and politics. Values in the Wild (Huang et al., 2025) analyzes over 300K human–AI conversations, building a taxonomy of 3,000 values and showing that in controversial situations, such as historical conflicts, Claude models often emphasize values of accuracy, human agency, and human wellbeing. These studies provide evidence that LLMs are being used in contexts with moral ambiguity.
Belief and Value Consistency. Many studies highlight inconsistency in LLMs. (Betz and Richardson, 2022) conducted a controlled study showing that LLMs produce logically inconsistent judgments (e.g., contradictory geographical relations or moral evaluations) even when trained on data from authors who individually maintain consistent beliefs. The MoralChoice benchmark (Scherrer et al., 2023) includes a subset of hypothetical high-ambiguity moral dilemmas, and results show that LLMs are indecisive within those contexts, indicating value indifference. The ValueConsistency benchmark (Moore et al., 2024) studies value consistency in controversial scenarios by paraphrasing prompts or translating them into different languages, then measuring whether the model’s binary stance (Yes / No) changes. Interestingly, many models outperform humans on paraphrase consistency. VAL-Bench differs by using prompt pairs that express opposing positions rather than paraphrases. We identify propositional belief contradictions in descriptive responses, using both the stated propositional attitude and its reasoning as validation. Our dataset is grounded in real-world issues, and has a rich diversity of contexts.
@@ 3 Dataset
We used the English content of Wikipedia (Wikimedia Enterprise (2024) to construct the VAL-Bench prompts dataset. We followed these steps to generate 114,745 pairs of prompts:
@@
    We filtered the Wikipedia sections using common terms used to designate controversial sections in Wikipedia, such as “Criticism” and “Scandals”.
@@
    We used LLM annotations to filter out sections that didn’t represent a divergent issue (due to errors in the heuristic used to find controversial sections).
@@
    We further prompted Gemma-3-27B-it (Team, 2025) to extract the pair of oppositely framed prompts from the text of a single controversial section in Wikipedia, each starting with Explain why. Each pair of prompts is referred to as for prompt and against prompt; there is no moral direction attached to this naming.
@@
Table 1: Top ten categories of issues with examples of issues in each category, illustrating the diversity of the dataset
Category % Example issues
Social and Cultural Issues 12.03 Racial bias in medical treatments, Machu Picchu artifacts in Yale University’s collection before 2011
Legal Disputes 5.55 Johnson v. Monsanto Co. over Roundup, Sexual allegations against the musician Basshunter
Religion 5.50 Child sexual abuse in Church, Ordaining women as Rabbis
History 4.75 Slavery in "Dutch Golden Age", Legacy of Russian Tsar Peter the Great
Media and Entertainment 3.83 Consolidation of news media in New Brunswick, Staged events in a Nature documentary
3.1 Dataset quality
Our evaluation strategy requires that the paired prompts in VAL-Bench have a common context and state contradictory positions. Human annotators were asked to label each pair of prompts on the dimensions of Common Context and Contradictory Position with binary True/False labels; a pair of prompts is well-formed if both are True. Two annotators independently labeled a set 500 pairs of prompts. The average well-formedness rating was 95.5% (Annotator 1: 91.8%, Annotator 2: 99.2%). Agreement between annotators was 91.8%. More details about this annotation process is provided in the Appendix.
In another annotation exercise, annotators noticed that some prompts conflate an actor relevant to the issue with the agent expected to answer the prompt (e.g., “Explain why you believe the allegations against you are false…”). This was reported twice in a set of 200 pairs of prompts.
@@ 3.2 Description of issues
Table 2: Top ten values encoded implicitly in the prompts. The dataset spans a wide spectrum of moral leanings.
Value % (count) Example issues
status quo reinforcement 3.08 (708) Opposition to Equal Rights Amendment in US, Re-interpretation of artistic themes in a stage play
We analyze the type of issues in the dataset across the dimensions of category and encoded values. [^6] LLMs were used for annotation, and their accuracy was not systematically validated. This analysis is presented primarily for illustrating diversity. The categories are shown in Table 1, listing 10 of the 20 categories used to classify each issue using an LLM-aided annotation process (the rest are listed in the Appendix).
The prompts implicitly encode specific values, as seen in the Figure 1. To create a taxonomy and understand which values are most commonly encoded, we use an LLM to annotate 22,950 test-set prompts with over 34,000 unique value strings, then condense them to 1,000 value strings using K-means clustering applied to embeddings. The top ten value strings are shown in Table 2 along with their frequency. Given the controversial nature of prompts, we expect to see morally good (accountability, prioritizing justice), morally bad (evading accountability), or morally ambiguous (tradition prioritization, legal formalism) values in that list. The prompts may also encode amoral values; for example, arguing for a traditional view on interpretation of art encodes the value of status quo reinforcement and tradition prioritization.
@@ 4 Evaluation
We formulate the process of evaluation as follows. The benchmark dataset ${\mathbb{D}}=\{(c_{i},p_{i}^{+},p_{i}^{-})\}_{i=1}^{N}$ consists of $N$ issues described with an issue description $c_{i}$ and paired prompts representing the for prompt and against prompt. We use an LM $\mathcal{M}$ to generate the response dataset ${\mathbb{G}}_{\mathcal{M}}=\{\mathcal{M}(p_{i}^{+}),\mathcal{M}(p_{i}^{-})\}_{i=1}^{N}$. To measure consistency, value preference and other metrics, we use an evaluator $\mathcal{E}$:
$\mathcal{E}:(c_{i},p_{i}^{+},r_{i}^{+},p_{i}^{-},r_{i}^{-})$ $\mapsto\bigl(\sigma_{i},\omega_{i},\psi(r_{i}^{+}),\nu(r_{i}^{+}),\psi(r_{i}^{-}),\nu(r_{i}^{-})\bigr)$ (1)
@@
Here, $\sigma_{i}=-2$ implies complete disagreement and $\sigma_{i}=2$ implies complete agreement between generated responses. $\omega_{i}=0$ denotes value indifference and $\omega_{i}=1$ denotes value preference. The PAC scores are adjusted as shown below:
$s_{i}=\begin{cases}\sigma_{i},&\text{if }\psi(r_{i}^{+})=0\text{ and }\psi(r_{i}^{-})=0,\\[6.0pt] max(1,\sigma_{i}),&\text{if }\psi(r_{i}^{+})+\psi(r_{i}^{-})=1,\\[6.0pt] 2,&\text{if }\psi(r_{i}^{+})+\psi(r_{i}^{-})=2.\end{cases}$ (6)
The score definition in equation 6 recognizes that refusals (responses stating that they can’t provide an opinion or take a stance) help in establishing consistency. For pairs with no-information responses (responses stating that there is insufficient information), we use the evaluator computed score $\sigma_{i}$.
The LLM responses are generated using nucleus sampling with generation settings dependent on the type of model (chat or thinking). The system prompt is consistently set to “You are a socially aware and responsible assistant. Answer the question honestly in 1-2 paragraphs.” More details are provided in the Appendix.
Table 3: Definition and description of metric symbols used to form a comprehensive picture of consistency, alignment and tradeoff with expressivity. Arrows denote the preferred direction of performance.
Symbol Definition Description
PAC $\uparrow$ $\tfrac{100}{4N}\sum_{i=1}^{N}(\sigma_{i}+2)$ Mean Position Alignment Consistency %
VPREF $\tfrac{100}{N}\sum_{i=1}^{N}\omega_{i}$ Mean Value Preference Rate %